# Offline Dedicated Quantum Attacks on Block Ciphers Based on Two Parallel Permutation-Based Pseudorandom Functions

### v2

@@ 4 Conclusion
This work successfully analyzed the quantum security of more instances of XOR-type function based on PolyMAC and TPP-PRFs, thereby resolving the open question raised by Shi et al. [23]. The dedicated quantum attack on BBB MACs required the adversary to perform superposition queries. We introduce a decoupled XOR-type function, termed $p$-XOR-type function, and propose an offline dedicated quantum attack on block ciphers based on TPP-PRFs. This attack yields significant efficiency improvements. In the Q2 model, the complexity of quantum query is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t})$ with the same time complexity. In the Q1 model, the offline attack can recover the key of TPP-PRFs, i.e., $l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$, by making $O(2^{p+t})$ classical queries and performing $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline computation time. It features a quantum-time/classical-query tradeoff of $T^{2}D=2^{2n-t}$, requiring $T=D=O(2^{(2n-t)/3})$, where $t$ is a tunable truncation parameter. Finally, we provide new insights into the quantum security of cryptographic schemes, including TPP-PRFs and specific instantiations such as $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
@@
Our results open several promising avenues for future research: First, explore the quantum security of more cryptographic structures that can construct decoupled XOR-type function in the Q1 model. Second, develop more generic and effective offline dedicated quantum attacks by identifying algebraic properties that apply to assess the quantum security of block ciphers, including $\mathsf{Feistel}$ [9, 14, 30], $\mathsf{MISTY}$ [6], $\mathsf{Lai\text{-}Massey}$ structure [20], and other permutation-based pseudorandom functions (e.g., $\mathsf{EDM}$ [8] and $\mathsf{EDMD}$ structure [21]).
@@ Acknowledgments
This work is supported by National Natural Science Foundation of China (Grant Nos. 62372048, 62272056, 62371069 )

### v3

@@ IV Conclusion
This work successfully analyzed the quantum security of PolyMAC and block cipher structures based on TPP-PRFs by constructing XOR-type functions, thereby resolving the open question raised by Shi et al. [shi2024dedicated]. We introduce a decoupled XOR-type function, termed $p$-XOR-type function, and propose an offline dedicated quantum attack on constructions based on TPP-PRFs. Our offline attack yields significant improvements. In the Q2 model, the quantum query complexity is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ with the same time complexity. In the Q1 model, our attack recovers the key of TPP-PRFs by making $O(2^{p+t})$ classical queries and performing $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline quantum computation time, which balances at $\tilde{O}(2^{(2n-t)/3})$. Our results provide new insights into the quantum security of cryptographic schemes, including TPP-PRFs and specific instantiations such as $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
@@
Our results open several promising avenues for future research: First, exploring the quantum security of more cryptographic structures that can construct decoupled XOR-type functions in the Q1 model. Second, developing more generic and effective offline dedicated quantum attacks by identifying algebraic properties applicable to attacking on block ciphers, including $\mathsf{Feistel}$ [dong2018quantum, ito2019quantum, sun2023quantumAQT], $\mathsf{MISTY}$ [canale2022simon], $\mathsf{Lai\text{-}Massey}$ structure [mao2022quantum], and other permutation-based pseudorandom functions (e.g., $\mathsf{EDM}$ [cogliati2016ewcdm] and $\mathsf{EDMD}$ structure [mennink2017encrypted]).
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant Nos. 62372048, 62272056, 62371069).